OpenStudy (anonymous):

(8x2 + 17x) + 6x2 = find the sum

OpenStudy (anonymous):

what are the like terms

OpenStudy (anonymous):

x?

OpenStudy (hoblos):

add like terms.. (8x2 + 17x) + 6x2 = 8x2 + 17x + 6x2 = (8+6)x2 +17x = 14x2 +17x

OpenStudy (anonymous):

@hoblos thanks

OpenStudy (hoblos):

my pleasure :)